A reservation clerk worked 9 hours one day. She spent twice as much time entering new
reservations as she did verifying old ones and one and a half as much time calling to confirm
reservations as verifying old ones. How much time did she spend entering new reservations?

Answer by mananth(16946)

verifying old ones.--------x hours
new reservations -----------------2x hours
calling to confirm reservations----3/2x
Total 9 hours
...
x+2x+3x/2 = 9
(2x+4x+3x) / 2 =9
9x =2*9
/9
x= 2 hours verifying old ones
2x = 2*2 = 4 hours ----new reservations
3x/2 = 3*2/2 = 3 hours calling to confirm.
...
mananth@hotmail.ca